OpenStudy (anonymous):

3 = 0.03t^3 + 0.05t^2 Solve for t

OpenStudy (anonymous):

Would I start by taking out t then moving 3 to the other side \[t(0.03t^2 + 0.05t) -3 =0\]

OpenStudy (mathmath333):

@aum

OpenStudy (aum):

3 = 0.03t^3 + 0.05t^2 0.03t^3 + 0.05t^2 - 3 = 0 Multiply by 100 to get rid of the decimals: 3t^3 + 5t^2 - 300 = 0 The LHS is not factorable. But the equation can be solved for t by graphing, by numerical analysis, etc. Which method do they want you to use?

OpenStudy (anonymous):

They do not specify any method but I guess graphing. I did by graphing and got that t= ~4.147

OpenStudy (aum):

Yes, t= ~4.147 is correct. If graphing is allowed then that would be the way to go.